# Progressive Growing of Video Tokenizers for Highly Compressed Latent Spaces

### v1

@@ Appendix D Resolving Jump Issues: Overlapping Frame Reconstruction and T2V
Appendix E CMD [43] Baseline Implementation Details.
Since the model weights for CMD [43]are not available in their official repository, we use the provided implementation and train with the originally provided settings to train on our internal data. We train on 256x256 resolution videos (because it was the maximum we could fit on A100 GPUs). We use a batch size of 1 per GPU but use 32 A100 GPUs to accommodate the recommended batch size by the authors for training. We also contacted the authors of CMD [43] for recommendations. The author recommended we also experiment with using the patch size of 8 (p=8) for potentially better efficiency and quality. Thus we provide results for both the settings in Table 1. We evaluate on a resolution of $512\times 512$. CMD transforms a video of 8 frames of $512\times 512$ into a content code $z_{c}$ of dimension $3\times 512\times 512$ and a motion code $z_{m}$ of dimension $8\times 8\times 512$. We found that CMD could only operate on square patches and not on a resolution of $360\times 640$.
@@
input size $17\times 256\times 256$ $17\times 256\times 256$ $17\times 256\times 256$
@@
latent dimension $5\times 32\times 32$ $3\times 32\times 32$ $2\times 32\times 32$
@@
input (latent) dimension $5\times 32\times 32$ $3\times 32\times 32$ $2\times 32\times 32$
output size $17\times 256\times 256$ $17\times 256\times 256$ $17\times 256\times 256$
@@
input size $17\times 256\times 256$
@@
kernel size $3\times 4\times 4$

### v2

@@ Appendix D Resolving Jump Issues: Overlapping Frame Reconstruction and Text-to-Video Generation
Appendix E Training Efficiency of Our Approach
Method Compression Time/iter Total iters Training time
(cumulative) (cumulative)
MagViT-v2 8×8×48$\times$ 8$\times$ 48 × 8 × 4 0.42s 300K 1.45 days
ProMAG 0.42s 300K 1.45 days
MagViT-v2 8$\times$8$\times$8 0.92s 400K 4.25 days
ProMAG (4$\times$ $\to$ 8$\times$) 0.42s 400K 1.94 days
MagViT-v2 8×8×168$\times$ 8$\times$ 168 × 8 × 16 1.04s 500K 6.01 days
ProMAG (8$\times$ $\to$ 16$\times$) 0.23s 500K 2.21 days
Table 6: Details of the per-iteration training time and total number of iterations comparing our method of progressive-growing (ProMAG) w.r.t to full model training (MagViT-v2). All the model has latent dimension ($z\_dim$) = 8. The training times are computed for 17 frames of 256$\times$256 videos on H100 GPU.
Our progressive-growing method does not cause any training efficiency overhead. In contrast, it improves the training efficiency compared to training the full model directly for high compression. From 6, we see that full model training from scratch (MagViT-v2) for 8$\times$ compression requires 2.3s/iter while the progressive growing strategy takes 1.05s/iter. Similarly, training 16$\times$ model on top of 8$\times$ model takes 0.57s/iter, compared to 2.6s/iter for full model training at 16$\times$ compression. As a result, progressive growth allows us to train faster while achieving significantly better performance with the same number of training iterations compared to the full training approach. The main reason why our approach of progressive growing has significantly lower training time per iteration is because it does not need a discriminator to grow the model from 4$\times$ to 8$\times$ or from 8$\times$ to 16$\times$ temporal compression. Additionally, most of the encoder and decoder layers are frozen. For fairness of comparison, we also train the full-model training (MagViT-v2) baseline for the same total number of iterations for a given temporal compression. Even though the full model training takes $\sim$3$\times$ more training time for 16$\times$ temporal compression compared to our method ProMAG, our method can still achieve a lot higher quality reconstruction.
Appendix F Reconstruction time Number of Parameters
Method Compression Enc. Time Dec. Time # Params
MagViT-v2 8$\times$8$\times$8 0.42 s 0.47 s 793 M
ProMAG 0.61 s 0.47 s 794 M
MagViT-v2 8$\times$8$\times$16 0.44 s 0.83 s 984 M
ProMAG 0.73 s 0.83 s 986 M
Table 7: Details of the reconstruction time and number of parameters of our model ProMAG w.r.t the baseline MagViT-v2. All the model has latent dimension ($z\_dim$) = 8. The times are computed for 17 frames of 512$\times$512 videos on A100 GPU.
Method Enc. (# Params) Enc. (Time) Dec. (# Params) Dec. (Time)
CogVideoX 92.2M 0.14s 123.3M 0.32s
WF-VAE 84.5M 0.02s 232.4M 0.14s
Wan 53.3M 0.07s 73.2M 0.12s
Hunyuan 100.3M 0.13s 146.1M 0.26s
CV-VAE 69M 0.05s 112M 0.15s
ProMAG 105M 0.08s 231M 0.14s
Table 8: Details of the reconstruction time and number of parameters of encoder and decoder of our model ProMAG w.r.t the baselines which perform 4$\times$ temporal compression. All the model has latent dimension ($z\_dim$) = 16. The times are computed for encoding and decoding 17 frames of 512$\times$512 resolution videos on A100 GPU.
For fairness of comparison, our model ProMAG follows the same encoder and decoder configuration to our implementation of the baseline MagViT-v2 for respective compression ratios. The difference is the progressive training strategy and the residual encoding with AdaNorm layers. From Table 7, the decoding time of our model and baseline MagViT-v2 is also the same because it follows the design. Only the encoding time of our model is  1.5$\times$ that of baseline MagViT-v2 because we need to do two forward passes through the encoder, one with the full input and the other with temporally subsampled input (subsampled by a factor of 2). Additionally, our model ProMAG has almost the same number of parameters as MagViT-v2; the only slight increase is due to the learnable parameters in AdaNorm blocks. Even so, for both 8$\times$ and 16$\times$ temporal compression, our model ProMAG has much better reconstruction quality than MagViT-v2 trained directly for 8$\times$ or 16$\times$ temporal compression. In Table 8, we show the number of parameters for encoder and decoder along with the encoding and decoding times between our model and other baselines with 4$\times$ temporal compression.
Note: Our focus (or contribution) is not to have the fastest or the most efficient video tokenizer.
@@ Appendix F Reconstruction time Number of Parameters
input size 17×256×25617$\times$ 256$\times$ 25617 × 256 × 256 17×256×25617$\times$ 256$\times$ 25617 × 256 × 256 17×256×25617$\times$ 256$\times$ 25617 × 256 × 256
@@
latent dimension 5×32×325$\times$ 32$\times$ 325 × 32 × 32 3×32×323$\times$ 32$\times$ 323 × 32 × 32 2×32×322$\times$ 32$\times$ 322 × 32 × 32
@@
input (latent) dimension 5×32×325$\times$ 32$\times$ 325 × 32 × 32 3×32×323$\times$ 32$\times$ 323 × 32 × 32 2×32×322$\times$ 32$\times$ 322 × 32 × 32
output size 17×256×25617$\times$ 256$\times$ 25617 × 256 × 256 17×256×25617$\times$ 256$\times$ 25617 × 256 × 256 17×256×25617$\times$ 256$\times$ 25617 × 256 × 256
@@
input size 17×256×25617$\times$ 256$\times$ 25617 × 256 × 256
@@
kernel size 3×4×43$\times$ 4$\times$ 43 × 4 × 4